# DreamAudio: Customized Text-to-Audio Generation with Diffusion Models

### v1

@@ VI-D2 Effectiveness of MRC
[…]
@@
VI-D3 Effectiveness of Data Masking and Dropping
We also conducted experiments on reference concept masking, dropping, and the impact of different data proportions. As shown in Table VI, the masking and dropping ratios significantly influence the model’s performance on customized tasks. When provided with more precise reference concepts, the model learns more effectively. However, based on the KL divergence and CLAPA scores, we observe that the generated results contain more similar information to the target audio. This may be because the model is led to pay attention to extraction and concatenation rather than generation, which may explain the reason of performance drop in general TTA generation tasks. This indicates that without masking and dropping, the model’s performance is degraded. Furthermore, in the data distribution comparison, we found that an excessive amount of customized-concatenation data makes the model overly reliant on the content of the reference concept. In contrast, Customized-Overlay data helps maintain a balance between leveraging the reference concept and the text prompts. Training with AudioCaps data improves the robustness of the model in general TTA scenarios. However, an excessive amount of AudioCaps data causes the model to shift focus away from the features provided by the reference concept, leading to degraded performance on the CTTA task.
@@ VI-E Limitation
Despite giving the SOTA performance, our method has certain limitations.
@@ VI-E1 Fixed Reference Format
VI-E2 The Number of References
The current model architecture limits the number of reference concept pairs to three, restricting the system’s flexibility when handling extensive or diverse customization. Ablation studies show that DreamAudio can be easily fine-tuned to handle a varying number of reference concepts. However, the performance of the system is dependent on the availability of appropriate training data.
VI-E3 Artificial Training Data
Due to the lack of real-world data for customized tasks, the audio generated by the proposed model can sometimes sound unnatural. Addressing this limitation requires the development of more diverse and high-quality datasets tailored to specific tasks, allowing models to capture nuanced patterns and improve their generalization to practical applications.
@@ VII Conclusion and Future Works
In this paper, we have introduced DreamAudio, an audio generation model for customized text-to-audio generation (CTTA). Given target prompts and reference audio-caption pairs as input, DreamAudio demonstrates SOTA performance in generating audio clips with specified audio contents in the CTTA task, while also delivering competitive results in the general TTA task, establishing a robust foundation for a wide range of audio-related applications. To further validate the real-world applicability of our method, we created the first benchmark for CTTA, which includes zero-shot audio events to simulate real-world scenarios. Our evaluation demonstrates that DreamAudio captures these user-specific audio events, highlighting its ability to generate meaningful and semantically coherent content. In addition, we studied the adaptability of the proposed method, which can effectively handle a varying number of audio-caption pairs as reference concepts. Future work will focus on improving the flexibility of incorporating reference audio-text pairs, changing the number of references, and developing diverse and high-quality real-world data. We will also extend the model to the integration of additional modalities such as images or video, and its use in other customized audio tasks, such as audio separation, style transfer, and editing.
@@ Acknowledgments
This research was partly supported by a research scholarship from the China Scholarship Council (CSC), an internship from ByteDance, funded by British Broadcasting Corporation Research and Development (BBC R&D), Engineering and Physical Sciences Research Council (EPSRC) Grant EP/T019751/1 “AI for Sound”, and a PhD scholarship from the Centre for Vision, Speech and Signal Processing (CVSSP), University of Surrey. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) license to any Author Accepted Manuscript version arising.

### v2

@@ VI-C Customized Generation on Various Number of References
AC $C$ CO Masking Dropping FAD $\downarrow$ KL $\downarrow$ CLAP $\uparrow$ CLAPA $\uparrow$ FAD $\downarrow$ KL $\downarrow$ CLAP $\uparrow$ CLAPA $\uparrow$
@@
VI-D Ablation Studies
VI-D1 Effectiveness of RFM
As shown in Table III, we compare the outputs of DreamAudio with different training approaches on various datasets. Experiments on RFM-based systems present enhanced performance on customized tasks, i.e., Customized-Concatenation and Customized-Overlay. On the other hand, we observe that DreamAudio-DDPM trained with diffusion-based techniques offers better results in the semantic feature than DreamAudio by achieving better CLAP scores. These experimental results demonstrate two key findings: a) RFM-based models show promising performance in generating customized and personalized content; b) DDPM-based approaches guide the models toward more robust performance by preserving the capability of generating features based on semantic information.
VI-D2 Effectiveness of UNet Backbone
We also compare our UNet architecture with diffusion-transformer (DiT) models, as DiT-based backbones have been shown to achieve superior performance in large-scale diffusion systems [55]. As reported in Table III, under a similar number of trainable parameters, DreamAudio-DiT-B does not achieve better performance than the UNet-based DreamAudio. By increasing the model size to approximately 1.31.3B, DreamAudio-DiT-L provides better results, however, it results in significantly longer training and inference time. These observations are consistent with previous findings that DiT architectures tend to outperform UNet models only at considerably larger scales (e.g., DiT-XL) [55]. Based on the current results, we adopt the UNet backbone in DreamAudio as a practical trade-off between performance and computational efficiency.
VI-D3 Effectiveness of MRC
We conducted various experiments with and without the external feature extraction path from the MRC module in Table III. For DreamAudio-UNet, the input of the generator module is a concatenation of general latent input and three VAE latent features of the reference audios. The results show that DreamAudio with only basic UNet structures exhibits significant performance degradation in all metrics. In addition, the comparison between DreamAudio-SEncoder and DreamAudio indicates that the use of two groups of encoder blocks for encoding and feature extraction helps improve the overall performance.
VI-D4 Effectiveness of Data Masking and Dropping
We also conducted experiments on reference concept masking, dropping, and the impact of different data proportions. As shown in Table VI, the masking and dropping ratios significantly influence the model’s performance on customized tasks. When provided with more precise reference concepts, the model learns more effectively. However, based on the KL divergence and CLAPA scores, we observe that the generated results contain more similar information to the target audio. This may be because the model is led to pay more attention to extraction and concatenation rather than generation, which may explain the reason of performance drop in general TTA generation tasks. This indicates that without masking and dropping, the model’s performance is degraded. Furthermore, in the data distribution comparison, we found that an excessive amount of customized-concatenation data makes the model overly reliant on the content of the reference concept. In contrast, Customized-Overlay data helps maintain a balance between leveraging the reference concept and the text prompts.
Training with AudioCaps data improves the robustness of the model in general TTA scenarios. However, an excessive amount of AudioCaps data causes the model to shift focus away from the features provided by the reference concept, leading to degraded performance on the CTTA task. Lastly, the experiment with 100% dropping represents the system trained without any reference text concepts while maintaining access to the reference audio clips, evaluating the ability of the model to perform customization based solely on the reference audio clips. The result shows a decrease in customization performance, which illustrates the importance of reference text in providing semantic grounding and helping the model exploit the reference audio features. In addition, as shown in Table IV, DreamAudio offers improved performance on general text-to-audio tasks on the AudioCaps benchmark, giving lower FAD and higher CLAP scores, as compared to AudioLDM and Stable Audio Open. This suggests that without the constraints of the reference text, DreamAudio tends to reduce its reliance on reference concepts and learn to generate audio based primarily on the target prompt.
@@ VI-E Limitation
Despite giving the satisfying performance, our method still has certain limitations.
@@ VI-E1 Fixed Reference Format
VI-E2 Fixed Audio Length
Our current architecture is primarily developed for 10-second audio, aligned with those in training data. Although the model can be extended to generate longer clips (e.g., 30 seconds or more) by adjusting the latent feature dimensions, we observe a noticeable performance degradation when samples significantly exceed the 10-second training window. To improve the model’s generalization across varying audio lengths, incorporating more diverse datasets with variable durations for training and evaluation remains an important aspect of future work.
VI-E3 The Number of References
The current model architecture limits the number of reference concept pairs to three, restricting the system’s flexibility when handling extensive or diverse customization. Ablation studies show that DreamAudio can be easily fine-tuned to handle a varying number of reference concepts. However, the performance of the system can be affected by the availability of appropriate training data.
VI-E4 Artificial Training Data
Due to the lack of real-world data for customized tasks, the audio generated by the proposed model can sometimes sound unnatural. In addition, in many cases, the reference audio naturally corresponds to a subset of the target audio events, which can make the task appear closer to stitching or inpainting in terms of the sound generated. Addressing this limitation requires the development of more diverse and high-quality datasets tailored to specific tasks, allowing models to capture detailed patterns and improve their generalization to practical applications.
@@ VII Conclusion and Future Works
In this paper, we have introduced DreamAudio, a model for customised text-to-audio generation (CTTA). Given target prompts and reference audio-caption pairs as input, DreamAudio demonstrates satisfying performance in generating audio clips with specified audio content in the CTTA task. In addition, DreamAudio delivers competitive results in the general TTA task, establishing a robust foundation for a wide range of audio-related applications. To further validate the real-world applicability of our method, we created the first benchmark for CTTA, which includes zero-shot audio events to simulate real-world scenarios. Our evaluation demonstrates that DreamAudio captures these user-specific audio events, highlighting its ability to generate meaningful and semantically coherent content. In addition, we studied the adaptability of the proposed method, which can effectively handle a varying number of audio-caption pairs as reference concepts. Future work will focus on improving the flexibility of incorporating reference audio-text pairs, changing the number of references, and developing diverse and high-quality real-world data. We will integrate additional modalities such as images and video, and extend its use in other customized audio tasks, such as audio separation, style transfer, and editing.
@@ Acknowledgments
This research was partly supported by a research scholarship from the China Scholarship Council (CSC), an internship from ByteDance, funding from British Broadcasting Corporation Research and Development (BBC R&D), Research England “Games and Innovation Nexus”, Engineering and Physical Sciences Research Council (EPSRC) Grant EP/T019751/1 “AI for Sound”, and a PhD scholarship from the Centre for Vision, Speech and Signal Processing (CVSSP), University of Surrey. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) license to any Author Accepted Manuscript version arising. The authors wish to thank the associate editor and the reviewers for their helpful comments to further improve this work.